If c(x)=4x-2 and d(x)=x^2+5x, what is c(d(x))?
Please help! I do not understand any of this!

Answer by jim_thompson5910(35256)

c(x)=4x-2

c(x)=4(x)-2

c(x)=4(x)-2

c(d(x))=4(d(x))-2 ... replace every x with d(x)

c(d(x))=4(x^2+5x)-2 ... replace the d(x) on the right side with x^2+5x

c(d(x))=4*x^2+4*5x-2

c(d(x))=4x^2+20x-2